The president of Big Fish GamesBig Fish Games​, an online gaming​ company, is considering the purchase of some equipment used for the development of new games. The cost is $400,000​,

the economic life and the recovery period are both 55 years, and there is no terminal disposal value. Annual pretax cash inflows from operations would increase by $130,000​, giving a total 55​-year pretax savings of $650,000 The income tax rate is 40​%, and the required​ after-tax rate of return is 14​%.

Compute the​ NPV, assuming​ straight-line depreciation of $80,000 yearly for tax purposes. Should Big Fish GamesBig Fish Games  acquire the​ equipment?

2.

Suppose the asset will be fully depreciated at the end of year 55 but can be sold for $25,000 cash. Should . Big Fish GamesBig Fish Games acquire the​ equipment? Show computations.

3.

Ignore number 2. Suppose the required​ after-tax rate of return is 10% instead of 14%. Should Big Fish GamesBig Fish Games acquire the​ equipment? Show computations.

Requirement 1. Compute the​ NPV, assuming​ straight-line depreciation of $80,000 yearly for tax purposes. Should Big Fish GamesBig Fish Games acquire the​ equipment?

Begin by computing the net present value​ (NPV) of the equipment investment. ​(Round dollar amounts the nearest whole number. Use a minus sign or parentheses for a negative net present value. Enter the present value factor to four decimal​ places, "X.XXXX.")

Present Value of

Ordinary Annuity of $1

Annual Cash

Total Present

at 5 years, 14%

Inflow

Value

Net present value:

Present value of annuity of equal annual:

After-tax cash flows from operations

x

per year =

After-tax cash savings from depreciation

x

per year =

Less: Initial investment

Net present value

Should Big Fish GamesBig Fish Games acquire the​ equipment?

Big Fish GamesBig Fish Games_____________ acquire the equipment because the NPV of the investment is ___________

Requirement 2. Suppose the asset will be fully depreciated at the end of year 55 but can be sold for $25,000 cash. Should Big Fish GamesBig Fish Games acquire the​ equipment? Show computations.

Begin by calculating the net present value of the equipment in this scenario. ​(Round dollar amounts the nearest whole number. Use a minus sign or parentheses for a negative net present value. Enter the present value factor to four decimal​ places, "X.XXXX.")

Present Value

of $1

Cash

Total Present

at 5 years, 14%

Inflow

Value

NPV of investment from requirement 1

Present value of lump sum after-tax cash flow

from disposal of equipment at the end of year 5

x

=

Net present value

Should Big Fish GamesBig Fish Games acquire the​ equipment? Big Fish GamesBig Fish Games______________ acquire the equipment because the NPV of the investment under this scenario is________.

Requirement 3. Ignore number 2. Suppose the required​ after-tax rate of return is 10%

instead of 14%. Should Big Fish GamesBig Fish Games acquire the​ equipment? Show computations.

Begin by calculating the net present value of the equipment in this scenario. ​(Round dollar amounts the nearest whole number. Use a minus sign or parentheses for a negative net present value. Enter the present value factor to four decimal​ places, "X.XXXX.")

Present Value of

Ordinary Annuity of $1

Annual Cash

Total Present

at 5 years, 10%

Inflow

Value

Net present value:

Present value of annuity of equal annual:

After-tax cash flows from operations

x

per year =

After-tax cash savings from depreciation

x

per year =

Less: Initial investment

Net present value

Should Big Fish Games acquire the​ equipment? Big Fish GamesBig Fish Games _________ acquire the equipment because the NPV of the investment under this scenario is _________.
Answer #1

Answer 1:

After tax cash flows from operations = 130000 * (1 - 40%) = $78,000

After-tax cash savings from depreciation = 80000 * 40% = $32,000

Big Fish Games should not acquire the equipment because the NPV of the investment is negative.

Answer 2:

After-tax cash flow from disposal of equipment at the end of year 5 = $25,000 * (1 - 40%) = $15,000

Big Fish Games should not acquire the equipment because the NPV of the investment under this scenario is negative.

Answer 3:

Big Fish Games should acquire the equipment because the NPV of the investment under this scenario is positive.